# LLM-driven Multimodal Target Volume Contouring in Radiation Oncology

### v3

@@ Abstract
Target volume contouring for radiation therapy is considered significantly more challenging than the normal organ segmentation tasks as it necessitates the utilization of both image and text-based clinical information. Inspired by the recent advancement of large language models (LLMs) that can facilitate the integration of the textural information and images, here we present a novel LLM-driven multimodal AI, namely LLMSeg, that utilizes the clinical text information and is applicable to the challenging task of target volume contouring for radiation therapy, and validate it within the context of breast cancer radiation therapy target volume contouring. Using external validation and data-insufficient environments, which attributes highly conducive to real-world applications, we demonstrate that the proposed model exhibits markedly improved performance compared to conventional unimodal AI models, particularly exhibiting robust generalization performance and data-efficiency. To our best knowledge, this is the first LLM-driven multimodal AI model that integrates the clinical text information into target volume delineation for radiation oncology.
@@
Department of Radiology, Massachusetts General Hospital and Harvard Medical School, Boston, MA, USA
Department of Radiation Oncology, Yonsei College of Medicine, Seoul, Korea
@@
Correspondence should be addressed to J.S.Kim and J.C.Ye
(jinsung@yuhs.ac, jong.ye@kaist.ac.kr)
Co-first authors
@@ Introduction
Inspired by the multimodal integration capability of LLMs and needs for multimodal information for tumor target delineation, here we present a novel multimodal clinical target volume (CTV) delineation model by integrating clinical text information through the LLM for conditioning a segmentation model. Specifically, by leveraging the textual information from well-trained LLMs through simple prompt tuning, our cross-attention-based segmentation model has adeptly integrated text-based clinical information into the target volume contouring task. More specifically, as illustrated in Fig. 1(a), we introduce an interactive alignment framework which uses both self-attention and cross-attention mechanisms in a bidirectional manner (text-to-image and image-to-text features), by following the concept of promptable segmentation from Segment Anything Model (SAM) [16]. To further improve the quality of features, we implement this interactive alignment between all the skip-connected image encoder features with the LLM feature. These layer-wise multimodal features are then combined to jointly predict the target labels through the multimodal decoder. In this way, we ensure the image encoder to efficiently extract meaningful text-related representations and vice versa. Finally, to transfer the LLM’s knowledge while the entire network parameters are kept and achieve superior performance in various downstream tasks [17, 18, 19], we adapt the idea of light-weight learnable text prompts to fully leverage the great linguistic capability of the LLM within the proposed multimodal AI framework.
Figure 1: (a) Illustration comparing the concept between the traditional vision-only AI and the multimodal AI in the context of radiotherapy target volume delineation. (b) Quantitative comparison of CTV contouring performance in the Dice metric. (c) Qualitative assessment of each concept.
As a proof of the concept, the model was applied to a breast cancer target volume delineation task further expanded to prostate cancer cases. By utilizing a well-curated, large-scale dataset from three institutions for development and external validation, we verify its capability to integrate pivotal clinical textural information, such as tumor stage, surgery type, and laterality. Experimental results confirm that the model not only demonstrates a significantly enhanced target contouring performance compared to existing unimodal segmentation models but also exhibits behavior that contours targets in accordance with provided clinical information. Notably, the model exhibits superior performance enhancement on an external dataset and shows stable performance gains in data-insufficient settings, demonstrating generalizability and data-efficiency that is not only apt for the characteristics of medical domain data but also align well with the perspective of clinical experts.
@@ Results
We qualitatively compared two different approaches in Fig. 1(c). In general, CTV for breast cancer radiation therapy can be categorized into two primary types: one that involves treatment of the breast or chest wall alone, and the other that electively treats the regional lymph nodal area (including axillary, supraclavicular, and internal mammary lymph nodes (LNs)) in addition to the aforementioned areas, given the frequent metastasis of breast cancer to these regions. On the left side of Fig. 1(c), despite the ground truth label posing CTV on both the breast and regional LNs, the vision-only model only contours the breast alone. Moreover, as the vision-only model lacks information about the laterality of the breast that diagnosed as cancer, partial segmentation masks are observed on the opposite breast. In contrast, the multimodal model accurately contours the breast and regional LNs that need to be treated as CTV. On the right side of Fig. 1(c), despite early breast cancer case requiring treatment of the breast only, the vision-only model incorrectly includes the regional LN as CTV. Moreover, CTVs are extended to the opposite breast. On the other hands, the multimodal model that integrates the clinical text information accurately contours the requisite treatment areas, encompassing both the breast and the regional LNs, aligning with the ground truth.
@@
We further compared our methods with other diverse vision-only and multimodal methods in Table 1. To the best of our knowledge, we firstly introduced a clinical information-driven 3D target contouring framework, in which the given textual information is not actually visible as an actual object in the input image, compared to current state-of-the-art (SOTA) multimodal segmentation models [20, 21]. Therefore, we adapted two publicly available 2D text-driven multimodal segmentation frameworks as our baseline multimodal models [22, 23]. Furthermore, we conducted comparisons with two advanced visual backbones [24, 25] to justify our selection of the 3D residual U-Net as the visual backbone. In results, LISA [22], regarded as the SOTA model for vision-language segmentation, showed inferior performance. This could be attributed to 2D vision-language models’s limitation of only accepting 2D images as input, and its fixed visual backbone optimized for the natural image domain lacks adaptability for 3D radiotherapy target volume segmentation, even after training with clinical dataset. On the other hand ConTEXTual Net [23], capable of handling 3D inputs with slight modification, showed promising performance. Nevertheless our approach demonstrated the SOTA performance across all evaluation metrics in various validation settings.
Table 1: Comparison of 3D CTV delineation performance between vision-only and multimodal AIs.
$\checkmark$ indicates the model selection for each vision-only or multimodal AI in all experiments.
Dataset Metric Vision-only AI Multimodal AI
3D Residual U-Net [26] 3D UNETR [24] 3D SegMamba [25] LISA[22] ConTEXTualNet [23] LLMSeg (Ours)
$\checkmark$ $\checkmark$
Internal Test (N=307) Dice $\uparrow$ 0.807 (0.788-0.825) 0.699 (0.679-0.718) 0.592 (0.576-0.606) 0.549 (0.522-0.573) 0.819 (0.800-0.835) 0.829 (0.809-0.845)
IoU $\uparrow$ 0.698 (0.677-0.718) 0.559 (0.538-0.580) 0.433 (0.418-0.447) 0.405 (0.383-0.425) 0.715 (0.695-0.733) 0.730 (0.709-0.748)
HD-95 $\downarrow$ 6.674 (5.891-7.452) 14.857 (14.258-15.483) 18.408 (17.930-18.918) 4.606 (3.963-5.311) 4.540 (3.806-5.273) 3.386 (2.890-3.949)
External Test #1 (N=206) Dice $\uparrow$ 0.731 (0.707-0.755) 0.555 (0.523-0.587) 0.522 (0.508-0.535) 0.535 (0.503-0.565) 0.815 (0.798-0.832) 0.822 (0.805-0.836)
IoU $\uparrow$ 0.599 (0.575-0.622) 0.422 (0.392-0.451) 0.359 (0.347-0.370) 0.394 (0.367-0.420) 0.701 (0.680-0.721) 0.709 (0.689-0.727)
HD-95 $\downarrow$ 19.922 (18.611-21.189) 16.451 (15.725-17.152) 22.677 (21.794-23.599) 5.257 (4.427-6.067) 6.362 (5.084-7.712) 4.256 (3.471-5.176)
External Test #2 (N=205) Dice $\uparrow$ 0.444 (0.419-0.469) 0.638 (0.619-0.658) 0.565 (0.554-0.576) 0.421 (0.379-0.459) 0.826 (0.809-0.840) 0.844 (0.826-0.857)
IoU $\uparrow$ 0.302 (0.282-0.322) 0.484 (0.464-0.507) 0.399 (0.388-0.409) 0.308 (0.276-0.339) 0.715 (0.697-0.732) 0.740 (0.722-0.756)
HD-95 $\downarrow$ 33.339 (32.997-33.615) 16.434 (15.935-16.904) 17.154 (16.761-17.486) 8.710 (7.583-9.828) 5.179 (4.250-6.160) 3.004 (2.555-3.533)
modified for 3D segmentation.
@@
The assessment of the target volume should not be based on mere metric evaluations such as the Dice, but rather by appropriate clinical rationale. In the context of breast contouring, this involves considerations such as whether the target volume has been contoured on the treated side of the breast, the contouring performed on the breast or chest wall depending on the type of surgery (breast-conserving surgery or mastectomy), and whether the regional LNs has been included. Therefore, the appropriateness of target contouring should be evaluated by a board-certified radiation oncologist, ensuring a clinically relevant perspective in the assessment. To this end, five rubrics (laterality, surgery type, volume definition, coverage, integrity) were suggested by the board-certified radiation oncologists, to objectively and specifically evaluate the target volume with differentiated scoring reflecting their importance. Detailed descriptions of these rubrics are available in Supplementary Table S1 with Figure S1.
@@
Table 2: Expert evaluation of CTV delineation performance.
Rubric Vision-only AI Multimodal AI (LLMSeg)
Laterality Surgery Type Volume Definition Coverage Integrity Total Laterality Surgery Type Volume Definition Coverage Integrity Total
(1 point) (1 point) (1.5 point) (1 point) (0.5 point) (5 point) (1 point) (1 point) (1.5 point) (1 point) (0.5 point) (5 point)
Internal Test (N=307) 0.786 (0.738-0.833) 0.887 (0.854-0.918) 0.900 (0.844-0.959) 0.478 (0.436-0.518) 0.216 (0.190-0.243) 3.267 (3.139-3.385) 0.990 (0.977-1.000) 0.987 (0.975-0.995) 1.142 (1.092-1.188) 0.602 (0.562-0.641) 0.253 (0.223-0.280) 3.973 (3.887-4.059)
External Test #1 (N=206) 0.344 (0.279-0.412) 0.863 (0.821-0.899) 0.680 (0.615-0.748) 0.186 (0.149-0.226) 0.124 (0.093-0.154) 2.198 (2.056-2.346) 0.990 (0.976-1.000) 0.983 (0.963-0.998) 1.174 (1.105-1.243) 0.532 (0.480-0.581) 0.260 (0.226-0.294) 3.939 (3.821-4.059)
External Test #2 (N=205) 0.268 (0.213-0.332) 0.828 (0.782-0.874) 0.488 (0.418-0.554) 0.029 (0.015-0.047) 0.087 (0.062-0.111) 1.700 (1.567-1.832) 0.990 (0.975-1.000) 0.986 (0.970-0.998) 1.173 (1.111-1.233) 0.611 (0.562-0.663) 0.250 (0.215-0.287) 4.010 (3.889-4.116)
@@
Figure 2: Comparison of target contouring performance when trained with different amounts of training dataset size. (a) Quantitative comparison for adll the validation sets in the Dice metric. (b) Qualitative comparison for external validation #1.
We observed the performance of each concept in target volume contouring by progressively reducing the size of training dataset. As illustrated in Fig. 2(a), our multimodal model demonstrated its data efficiency by maintaining stable performance above 0.8 in the Dice even with 40% of data availability. This starkly contrasts with the vision-only model, whose performance dropped from initial Dice of 0.8 to 0.7. When we utilizing only 20% of the training dataset, the multimodal model’s performance decreased slightly below 0.8 in the Dice, while the vision-only model completely failed to contour CTV in the limited dataset scenario. This performance gap was particularly evident in the external validation results. For external validation #1, the initial discrepancy between two models was approximately 0.1 in the Dice metric. However, as the size of training dataset decreased, the discrepancy doubled. For external validation #2, notable overfitting issues were observed in the vision-only model. On the contrary, our multimodal model achieved robust performance when trained with a reduced dataset of less than 40%. Qualitative analysis, as depicted in Fig. 2(b), also supports these results. Detailed quantitative results for all metrics are further provided in Supplementary Table S3.
Figure 3: (a) Illustration of modification of the input textual information, given the same CT scan, and (b-c) example cases for radiotherapy target contouring.
@@
As depicted in Fig. 3(b)-(c), the model performed contouring different targets for the same CT, contingent on the provided textual information. In Fig. 3(b), for a patient with left breast cancer at stage T1N0M0, upstaging the T stage or N stage demonstrated the inclusion of regional LNs, and altering the tumor’s laterality from left to right resulted in contouring on the opposite side. Interestingly, when the type of surgery was changed from breast conserving surgery (BCS) to total mastectomy, it was observed that the previously spared skin was no longer spared, and the target volume was expanded to include the chest wall. For another patient with right breast cancer at stage T2N1M0, as exemplified in Fig. 3(c), downstaging N stage leads to the omission of regional LNs from the designated target volume, and changing the type of surgery to BCS results in a strategic shift to sparing the skin and excluding the pectoralis muscle from the treatment volume. These quantitative results align precisely with the decision policy of radiation oncologists, and substantiate that our model contours the target volume, strongly referencing the textual clinical information as well as the imaging features.

### v4

@@ Abstract
Target volume contouring for radiation therapy is considered significantly more challenging than the normal organ segmentation tasks as it necessitates the utilization of both image and text-based clinical information. Inspired by the recent advancement of large language models (LLMs) that can facilitate the integration of the textural information and images, here we present an LLM-driven multimodal artificial intelligence (AI), namely LLMSeg, that utilizes the clinical information and is applicable to the challenging task of 3-dimensional context-aware target volume delineation for radiation oncology. We validate our proposed LLMSeg within the context of breast cancer radiotherapy using external validation and data-insufficient environments, which attributes highly conducive to real-world applications. We demonstrate that the proposed multimodal LLMSeg exhibits markedly improved performance compared to conventional unimodal AI models, particularly exhibiting robust generalization performance and data-efficiency.
@@
Department of Radiology, Massachusetts General Hospital (MGH) and Harvard Medical School, Boston, MA, USA
Department of Radiation Oncology, Yonsei University College of Medicine, Seoul, Korea
@@
These authors contributed equally
Correspondence to Jin Sung Kim (jinsung@yuhs.ac) or Jong Chul Ye (jong.ye@kaist.ac.kr)
@@ Introduction
Inspired by the multimodal integration capability of LLMs and needs for multimodal information for tumor target delineation, here we present a 3-dimensional (3D) multimodal clinical target volume (CTV) delineation model, LLMSeg, by integrating clinical information through the LLM for conditioning a segmentation model. Specifically, by leveraging the textual information from well-trained LLMs through simple prompt tuning, our cross-attention-based segmentation model has adeptly integrated text-based clinical information into the target volume contouring task. More specifically, as illustrated in Fig. 1(a), we introduce an interactive alignment framework which uses both self-attention and cross-attention mechanisms in a bidirectional manner (text-to-image and image-to-text features), by following the concept of promptable segmentation from Segment Anything Model (SAM) [16]. To further improve the quality of features, we implement this interactive alignment between all the skip-connected image encoder features with the LLM feature. These layer-wise multimodal features are then combined to jointly predict the target labels through the multimodal decoder. In this way, we ensure the image encoder to efficiently extract meaningful text-related representations and vice versa. Finally, to transfer the LLM’s knowledge while the entire network parameters are kept and achieve superior performance in various downstream tasks [17, 18, 19], we adapt the idea of light-weight learnable text prompts to fully leverage the great linguistic capability of the LLM within the proposed multimodal AI framework.
In this work, we apply LLMSeg to breast cancer target volume delineation task to evaluate its context-aware radiotherapy target delineation performance compared to unimodal AI. Additionally, we expand its application to prostate cancer cases. By utilizing a well-curated, large-scale dataset from three institutions for development and external validation, we verify its capability to integrate pivotal clinical information, such as tumor stage, surgery type, and laterality. Experimental results confirm that the model not only demonstrates a significantly enhanced target contouring performance compared to existing unimodal segmentation models but also exhibits behavior that contours targets in accordance with provided clinical information. Notably, the model exhibits superior performance enhancement on an external dataset and shows stable performance gains in data-insufficient settings, demonstrating generalizability and data-efficiency that is not only apt for the characteristics of medical domain data but also align well with the perspective of clinical experts.
@@ Results
We qualitatively compared two different approaches in Fig. 1(c). In general, CTV for breast cancer radiation therapy can be categorized into two primary types: one that involves treatment of the breast or chest wall alone, and the other that electively treats the regional lymph nodal area (including axillary, supraclavicular, and internal mammary lymph nodes (LNs)) in addition to the aforementioned areas, given the frequent metastasis of breast cancer to these regions. On the left side of Fig. 1(c), despite the ground truth label posing CTV on both the breast and regional LNs, the vision-only model only contours the breast alone. Moreover, as the vision-only model lacks information about the laterality of the breast that diagnosed as cancer, partial segmentation masks are observed on the opposite breast. In contrast, the multimodal model accurately contours the breast and regional LNs that need to be treated as CTV. On the right side of Fig. 1(c), despite early breast cancer case requiring treatment of the breast only, the vision-only model incorrectly includes the regional LN as CTV. Moreover, CTVs are extended to the opposite breast. On the other hands, the multimodal model that integrates the clinical information accurately contours the requisite treatment areas, encompassing both the breast and the regional LNs, aligning with the ground truth.
@@
We further compared our methods with other diverse vision-only and multimodal methods in Table 1. Our proposed context-aware segmentation, in which the given textual information is not explicitly visible as an actual object in the input image, compared to traditional vision-language segmentation [20, 21]. Therefore, we adapted publicly available 2D text-driven multimodal segmentation frameworks from various segmentation categories as our baseline models [22, 23, 24]. Furthermore, we conducted comparisons with two advanced visual backbones [25, 26] to justify our selection of the 3D residual U-Net as the visual backbone. In the results shown in Table 1, HIPIE [22] and LISA [23], considered SOTA models for 2D referring and reasoning segmentation respectively, showed suboptimal performance in 3D context-aware segmentation. On the other hand ConTEXTualNet [24], capable of handling 3D images as inputs, showed promising performance. Nevertheless our approach demonstrated the SOTA performance across all evaluation metrics in various validation settings.
@@
The assessment of the target volume should not be based on mere metric evaluations such as the Dice, but rather by appropriate clinical rationale. In the context of breast contouring, this involves considerations such as whether the target volume has been contoured on the treated side of the breast, the contouring performed on the breast or chest wall depending on the type of surgery (breast-conserving surgery or mastectomy), and whether the regional LNs has been included. Therefore, the appropriateness of target contouring should be evaluated by a board-certified radiation oncologist, ensuring a clinically relevant perspective in the assessment. To this end, five rubrics (laterality, surgery type, volume definition, coverage, integrity) were suggested by the board-certified radiation oncologists, to objectively and specifically evaluate the target volume with differentiated scoring reflecting their importance. Detailed descriptions of these rubrics are available in Supplementary Table 1 with Supplementary Fig. 1.
@@
We observed the performance of each concept in target volume contouring by progressively reducing the size of training dataset. As illustrated in Fig. 2(a), our multimodal model demonstrated its data efficiency by maintaining stable performance above 0.8 in the Dice even with 40% of data availability. This starkly contrasts with the vision-only model, whose performance dropped from initial Dice of 0.8 to 0.7. When we utilizing only 20% of the training dataset, the multimodal model’s performance decreased slightly below 0.8 in the Dice, while the vision-only model completely failed to contour CTV in the limited dataset scenario. This performance gap was particularly evident in the external validation results. For external validation #1, the initial discrepancy between two models was approximately 0.1 in the Dice metric. However, as the size of training dataset decreased, the discrepancy doubled. For external validation #2, notable overfitting issues were observed in the vision-only model. On the contrary, our multimodal model achieved robust performance when trained with a reduced dataset of less than 40%. Qualitative analysis, as depicted in Fig. 2(b), also supports these results. Detailed quantitative results for all metrics are further provided in Supplementary Table 2.
@@
As depicted in Fig. 3(b)-(c), the model performed contouring different targets for the same CT, contingent on the provided clinical data. In Fig. 3(b), for a patient with left breast cancer at stage T1N0M0, upstaging the T stage or N stage demonstrated the inclusion of regional LNs, and altering the tumor’s laterality from left to right resulted in contouring on the opposite side. Interestingly, when the type of surgery was changed from breast conserving surgery (BCS) to total mastectomy, it was observed that the previously spared skin was no longer spared, and the target volume was expanded to include the chest wall. For another patient with right breast cancer at stage T2N1M0, as exemplified in Fig. 3(c), downstaging N stage leads to the omission of regional LNs from the designated target volume, and changing the type of surgery to BCS results in a strategic shift to sparing the skin and excluding the pectoralis muscle from the treatment volume. These quantitative results align precisely with the decision policy of radiation oncologists, and substantiate that our model contours the target volume, strongly referencing the textual clinical information as well as the imaging features.